Solve the equation:
4^(x) + 6^(x) = 9^(x)
1
Expert's answer
2021-01-14T12:50:53-0500

"4^x+6^x=9^x\\"

Will divide by "4^x"


"1+\frac{6^x}{4^x}=\frac{9^x}{4^x}\\1+(\frac{3}{2})^x=(\frac{3}{2})^{2x}"


Put "(\frac{3}{2})^x=t"


Then

"t^2-t-1=0"


"t=\frac{1 \mp \sqrt{1^2-4\times1\times(-1)}}{2\times1}\\t=\frac{1\mp \sqrt{5}}{2}\\"


t should have only positive value because 3/2 is positive value and power of positive number always give positive value.

Thus

"t=\frac{1+\sqrt{5}}{2}"


"\frac{1+\sqrt{5}}{2}=(\frac{3}{2})^x
\\and\\ln(\frac{1+\sqrt{5}}{2})=xln(\frac{3}{2})\\x=1.23"